Good afternoon and happt thanksgiving.
I am working on some math homework and could use a littlel help.
Here is the question.
A line passing through (5, –6) and (–9, y) is perpendicular to a line with slope 14/5 . Find the value of y.
Thanks you tutors rock

Answer by checkley71(8403)

ANY LINE PERPENDICULAR TO A LINE WITH A SLOPE OF 14/5 HAS A SLOPE OF -5/14.
THEN (Y2-Y1)/(X2-X1)=-5/14
(Y2+6)/(-9-5)=-5/14
(Y2+6)/-14=-5/14
-(Y2+6)14=-5/14 BECAUSE THE DENOMINATORS ARE THE SAME THE NUMERATORS ARE EQUAL.
-(Y2+6)=-5
-Y2-6=-5
-Y2=-5+6
-Y2=1
Y2=-1 ANSWER.
PROOF
(-1+6)/-14=-5/14
5/-14=-5/14
-5/14=-5/14